OpenStudy (bloomlocke367):

What are the vertices of the hyperbola with equation 4y^2-25x^2=100? Graph the hyperbola

OpenStudy (bloomlocke367):

@iGreen

OpenStudy (anonymous):

do you know what it looks like? you need to know that first

OpenStudy (bloomlocke367):

All I know is that it's a hyperbola

OpenStudy (anonymous):

which way it it oriented ?

OpenStudy (anonymous):

if you know that, we can answer the question instantly more or less

OpenStudy (anonymous):

hint: you can tall by which comes first, the x or the y

OpenStudy (anonymous):

*tell

OpenStudy (bloomlocke367):

I don't know... y comes first..

OpenStudy (anonymous):

yes, we see the y comes first that means it looks like the one on the right, not on the left

OpenStudy (bloomlocke367):

okay.

OpenStudy (bloomlocke367):

how do I find the vertices? do I need to divide by 100?

OpenStudy (anonymous):

divide all by 100 to put it in the standard form,, you get \[\frac{y^2}{25}-\frac{x^2}{16}=1\]